# Congruences modulo powers of 3 for 6-colored generalized Frobenius partitions

### v1

@@ 3. Modular equation
Given the form of $L_{1}$ in (3.6), we expect to represent each $L_{\alpha}$ in terms of $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n};$
@@
for $\alpha\geq 1$. Recalling that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
@@
For integer $\alpha\geq 1$, we have
@@
which demonstrates that (4.11) and (4.13) hold for $\alpha=1$. Next, we suppose (4.11) and (4.13) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
@@
By comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ in the equation above, we have
@@
Hence, for $n\geq 1$, we have
$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
@@
which indicates (4.12) holds for $\alpha=s$. For $d_{0}^{(2s)}$, from (4.15) we obtain that
@@
By the assumption, we have $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$. We denote $[t^{k}]f$ to the coefficients of $t^{k}$ in $f$. With (3.9), we have
@@
In the Appendix, we get that $b(-1,0,0)=12$, $b(-1,1,0)=36$, and $b(-1,2,0)=69$ which all can be divided by $3$. By the recurrence relation above, we attain that $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$. Thus, we get that (4.14) holds for $\alpha=s$.
Finally, we suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.4 and Lemma 4.6, we have
@@
Equating the coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ in the equation above, we have
@@
Hence, for $n\geq-1$, we have
$\pi(d_{n}^{(2s+1)})$ $\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
@@
which indicates (4.11) holds for $\alpha=s+1$. From (4.17), we have
@@
By the assumption, we have $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$. Similar to (4.16), we obtain that
@@
Hence, we attain that $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$ so that (4.13) holds for $\alpha=s+1$, which completes the proof of Lemma 4.7. ∎
@@ Acknowledgements
The first author was supported by the National Key R&D Program of China (Grant No. 2024YFA1014500). The first author was also supported in part by the National Natural Science Foundation of China (Grant No. 12201387).

### v2

@@ 3. Modular equation
Given the form of $L_{1}$ in (3.2), we expect to represent each $L_{\alpha}$ in the form $\frac{t^{n}}{(1-3t)^{m}}$ for $n\geq-1$ and $m\geq 0$. If we set
$L_{2\alpha-1}=p_{0}y^{3^{2\alpha}-1}\sum_{n\geq-1}d_{n}^{(2\alpha-1)}t^{n},$
@@
for $\alpha\geq 1$, and recall that $\pi(n)$ denotes the 3-adic order of n (i.e. the highest power of 3 that divides $n$), then Theorem 1.1 follows from the next lemma.
@@
For every integer $\alpha\geq 1$, we have
@@
which shows that $\pi(d_{n}^{(1)})\geq\lfloor\frac{2n+5}{3}\rfloor+1$ and $\pi(d_{0}^{(1)})=4$. Thus, (4.11) and (4.13) hold for $\alpha=1$.
Next, suppose (4.11) and (4.13) hold for $\alpha\leq s$. We want to use $U_{B}$ operator on $L_{2s-1}$ to get $L_{2s}$, especially the coefficients $d_{n}^{(2s)}$ for $n\geq 0$. Then by Definition 2.3 and Lemma 4.6, we have
@@
Comparing the coefficients of $p_{1}y^{3^{2s+1}-3}t^{n}$ on both sides, we have
@@
We notice that $\pi(n)$ has the properties that $\pi(x+y)\geq\min\{\pi(x),\pi(y)\}$ and $\pi(xy)=\pi(x)+\pi(y)$ for $x,y\in\mathbb{Z}$. According to (4.10) and (4.11), we obtain that $\pi(b(k,3^{2s}-1,n))\geq\lfloor\frac{2n-k+2}{3}\rfloor$ and $\pi(d_{k}^{(2s-1)})\geq\lfloor\frac{2k+5}{3}\rfloor+s$. Hence, for $n\geq 1$,
$\pi(d_{n}^{(2s)})$ $\geq\min_{k\geq-1}\{\pi(d_{k}^{(2s-1)})+\pi(b(k,3^{2s}-1,n))\}$
$\geq\min_{k\geq-1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+2}{3}\rfloor\}$
@@
which indicates that (4.12) holds for $\alpha=s$.
For $d_{0}^{(2s)}$, from (4.15), we obtain that
@@
By the induction hypothesis, $\pi(d_{-1}^{(2s-1)})\geq s+1$, $\pi(d_{0}^{(2s-1)})\geq s+2$, and $\pi(d_{1}^{(2s-1)})\geq s+2$.
Let $t^{k}]f$ denote the coefficients of $t^{k}$ in $f$. Using(3.5), we find
@@
From the Appendix, b(1,0,0)=12b(-1,0,0)=12, b(1,1,0)=36b(-1,1,0)=36, and b(1,2,0)=69b(-1,2,0)=69, all divisible by $3$. By the recurrence above, $3\mid b(-1,m,0)$ for all $m\in\mathbb{N}$. Hence, (4.14) holds for $\alpha=s$.
Now suppose (4.12) and (4.14) hold for $\alpha\leq s$. Then by Definition 2.3 and Lemma 4.6,
@@
Equating coefficients of $p_{0}y^{3^{2s+2}-1}t^{n}$ on both sides, we get
@@
According to (4.9), (4.12), and (4.14), we obtain that $\pi(a(k,3^{2s+1}-3,n))\geq\lfloor\frac{2n-k+3}{2}\rfloor$, $\pi(d_{k}^{(2s)})\geq\lfloor\frac{2n+5}{3}\rfloor+s$ for $k\geq 1$, and $\pi(d_{0}^{(2s)})\geq s+2$. Thus, for $n\geq-1$,
$\pi(d_{n}^{(2s+1)})$ $\geq\min_{k\geq 0}\{\pi(d_{k}^{(2s)})+\pi(a(k,3^{2s+1}-3,n))\}$
$\geq\min\{\min_{k\geq 1}\{\pi(d_{k}^{(2s)})+\pi(a(k,3^{2s+1}-3,n))\},\pi(d_{0}^{(2s)})+\pi(a(0,3^{2s+1}-3,n))\}$
$\geq\min\{\min_{k\geq 1}\{\lfloor\frac{2k+5}{3}\rfloor+s+\lfloor\frac{2n-k+3}{3}\rfloor\},s+2+\lfloor\frac{2n+3}{3}\rfloor\}$
$\geq\min\{\min\{\min_{k\geq 2}\{\lfloor\frac{2n+k+3}{3}\rfloor+s+1\},\lfloor\frac{2n+5}{3}\rfloor+s+1\},\lfloor\frac{2n+6}{3}\rfloor+s+1\}$
@@
where the forth inequality is obtained by splitting the part where $k\textgreater 1$ into two cases: $k=1$ and $k\geq 2$. Thus, (4.11) holds for $\alpha=s+1$.
From (4.17), we have
@@
By the induction hypothesis, $\pi(d_{0}^{(2s)})\geq s+2$, $\pi(d_{1}^{(2s)})\geq s+2$, $\pi(d_{2}^{(2s)})\geq s+3$, and $\pi(d_{3}^{(2s)})\geq s+3$.
Similar to (4.16), we derive
@@
Hence, $3\mid a(0,m,0)$ and $3\mid a(1,m,0)$ for all $m\in\mathbb{N}$, so (4.13) holds for $\alpha=s+1$. This completes the proof of Lemma 4.8. ∎
@@ Acknowledgements
The first author was supported by the National Key R&D Program of China (Grant No. 2024YFA1014500) and the National Natural Science Foundation of China (Grant No. 12201387).